if (a,1/a),(b,1/b)&(c,1/c)are the vertices of a triangle,(a,b) are roots of x2-6p1x+2=0,(b,c)roots of x2-6p2x+3=0&(c,a)roots of x2-6p3x+6=0.p1,p2,p3being positive,find p1,p2,p3¢roid of triangle.

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jagannath

(a,b) are roots of x2-6p1x+2=0

a+b=6p1  ............1

ab=2 .............2

(b,c)roots of x2-6p2x+3=0

b+c=6p2............3

bc=3..........4

(c,a)roots of x2-6p3x+6=0

c+a=6p3 .............5

ca=6 .............6

 

from equation 2,4,6

(ab)(bc)=ca

b=1

and a=2

c=3

so P1 =1/2

p2 =2/3

P3=5/6

 

now centroid ={(1+2+3)/3    ,(1/2 +1/1  +1/3 )/3 }

                          (2,11/18)


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin